Trace the curve (x²+y²)x=ay², a>0 stating all the properties used in the process
1
Expert's answer
2020-12-02T17:45:27-0500

"\displaystyle

\textbf{\textsf{Intercepts}}\\

(x^2+y^2)x=ay^2\\

\textsf{At}\,\, x = 0, \,y = 0\\

\therefore\textsf{The curve passses through}\\
\textsf{the origin.}\\

\textbf{\textsf{Stationary points}}\\

x^3 + y^2x = ay^2\\

3x^2 + y^2 + 2xyy' = 2ayy'\\


\frac{3x^2 + y^2}{2ay - 2xy} = y'\\

\textsf{The value of}\,x \, \textsf{for which the}\\
\textsf{there is is a stationary point}\, y' = 0\\

3x^2 = -y^2\\

y = j\sqrt{3}x\, \textsf{where}\, j = \sqrt{-1}\\


\textsf{For the line}\, y = j\sqrt{3}x\,, \textsf{for all values}\\
\textsf{of}\,x, \textsf{there are no}\\
\textsf{values of}\, y \, \textsf{that exists}\\
\textsf{in the real plane and vice-versa}\\
\textsf{except at}\, (x, y) = (0, 0). \\


\textsf{Therefore, there is no maximum value or minimum value}\\
\textsf{of}\,y\, \textsf{and there is a stationary point}\, x = 0. \, \\
\textsf{Since at}\, \frac{\mathrm{d}y}{\mathrm{d}x} = 0\,\, \textsf{at}\,\, x = 0.\\

\textbf{\textsf{Asymptotes}}\\

x^3 + y^2x = ay^2\\


x^3 = y^2(a - x)\\


y = \pm\sqrt{\frac{x^3}{a - x}}\\


\textsf{The line}\,\,x = a \, \, \textsf{is a vertical asymptote}\\


\textsf{And as}\,\, x \rightarrow a, y \rightarrow \pm\infty.\\

\textsf{Therefore, the graph is increasing}\\
\textsf{and decreasing for}\,\,
x > 0."